# DynEx: Dynamic Code Synthesis with Structured Design Exploration for Accelerated Exploratory Programming

### v1

@@ 4.4.3. Claude’s immediate feedback loop is engaging
However, participants did also note that this was a drawback in terms of enabling a thorough design space exploration – because of Claude Artifact’s inherently linear workflow, users only appended features to their initial prototype, as opposed to trying out different ideas. P4 noted: ”[Claude Artifact] just did what I told it to do… [the prototype] was what I expected”. P8 stated: ”I couldn’t even brainstorm the fact that it could brainstorm,” reflecting a broader trend among participant usage patterns with Claude Artifact; participants did not spend much time ideating with Claude Artifact because it was not intuitive on how they could do so.
@@ 5. Discussion
DynEx successfully enabled exploratory programming, facilitating the transformation of abstract ideas into concrete prototypes. We discuss how LLMs can significantly enhance and accelerate the process of exploratory programming. We additionally consider the implications of unstructured and structured approaches towards design exploration. Finally, we discuss the potential of the Design Matrix acting as a liaison between designers and engineers.
@@ 5.2. Structured vs Unstructured Design Exploration
While users felt supported by DynEx’s structured design space exploration, they also appreciated aspects of Claude’s chat-interface – namely, that it provided a more immediate feedback loop. Claude Artifact’s interactions are simple and direct – users could type an input and almost instantly receive a code output. This responsiveness engaged users, motivating them to iterate more quickly on ideas. Chatbot systems are very common; people keep making them, and users are very familiar with them. They invite open-ended input, allowing users be highly specific or intentionally vague in their prompts. However, just because it is common, does not mean it is the best paradigm. Comparing it effectiveness to a structured UI, such as DynEx, is also important.
@@
5.3. Bridging the gap between design and programming
DynEx brings together design and programming for the user. Traditionally, designers and engineers are siloed during software development - designers design and engineers implement, with product managers attempting to bridge the gap between the two. However, the design landscape is always changing, and context is bound to be lost – even a product manager may not be up to date on the latest changes. Addressing this disconnect between design and engineering is challenging but likely useful; our user study with individuals demonstrated the value in contextualizing implementations in design. However, in the working world, designers and engineers work in teams. Designers must continuously update engineers through shared Figma files as designs change. LLMs offer a unique opportunity to provide a continuous feedback loop between teams of designers and engineers and reduce this friction. Existing LLM-powered tools for established codebases, such as CoPilot and Gemini, only support developers in implementation. DynEx supported exploratory programming with design exploration for individuals with successful results; the idea of incorporating design context into these LLM-powered systems could extend to tools for software development workflow on teams.
@@ 6. Limitations and Future Work
Our user study was limited to ten participants, all of whom had computer science backgrounds and many years of technical experience. This may not be representative of the broader population that could benefit from an exploratory programming system – namely, anyone with an idea to prototype, regardless of their past technical exposure. Future studies should expand the demographics of participants to best inform the development of a system for users of varying technical backgrounds. Our study was also limited in the amount of time participants had to prototype. They had 30 minutes with each system – with additional time, participants may have created a variety of prototypes and compared them with one another to emulate a richer prototyping experience. Future studies should examine how users interact with the system when given more time.
Our user study also only used Claude Artifact as a baseline. Although we conducted a brief exploration which suggested Claude Artifact as being the most suitable as a baseline compared to other existing LLM-based tools, future work could involve comparisons between more systems in order to identify specific components or design-choices that are effective in supporting exploratory programming. Furthermore, while users were given an introduction to Claude Artifact, they were still quite inexperienced; experienced users might use Claude Artifact as a tool for prototyping differently than novices.
@@ 7. Conclusion
The advancement of LLMs poses a unique opportunity for exploratory programming. LLMs can generate code rapidly and also help users enhance and refine their problem space. In this paper, we presented DynEx, a system to accelerate exploratory programming through a structured, LLM-guided Design Matrix and implementation through iterative LLM code-generation. We also introduce self-invoking multi-modal LLMs as a technique for LLM-based exploratory programming, allowing for faster creation of diverse and functional prototypes. Through our user evaluation with 10 technical experts, we found that DynEx’s Design Matrix allowed to users to explore, refine, and realize feature-rich applications suitable for prototyping. We conclude by discussing how systems like DynEx can help bridge the gap between design and implementation, empowering a broad range of individuals to bring their ideas to life.
@@ Appendix A Few-shot Examples for the Design Matrix
Figure 8. Few-shot examples for entries in the Design Matrix
[…]
¡Few-shot Examples for the Design Matrix¿
[…]

### v3

@@ 6.4.3. Claude’s immediate feedback loop is engaging
However, participants did also note that this was a drawback in terms of enabling a thorough design space exploration — because of Claude Artifact’s inherently linear workflow, users only appended features to their initial prototype, as opposed to trying out different ideas. P4 noted: ”[Claude Artifact] just did what I told it to do… [the prototype] was what I expected”. P8 stated: ”I couldn’t even brainstorm the fact that it could brainstorm,” reflecting a broader trend among participant usage patterns with Claude Artifact; participants did not spend much time ideating with Claude Artifact because it was not intuitive on how they could do so.
@@ 7.2. Structured vs Unstructured Design Exploration
While users felt supported by DynEx’s structured design space exploration, they also appreciated aspects of Claude’s chat-interface — namely, that it provided a more immediate feedback loop. Claude Artifact’s interactions are simple and direct — users could type an input and almost instantly receive a code output. This responsiveness engaged users, motivating them to iterate more quickly on ideas. Chatbot systems are very common; people keep making them, and users are very familiar with them. They invite open-ended input, allowing users be highly specific or intentionally vague in their prompts. However, just because it is common, does not mean it is the best paradigm. Comparing it effectiveness to a structured UI, such as DynEx, is also important.
@@
7.3. Bridging the gap between design and engineering
DynEx brings together design and implementation through the Design Matrix and code generation. Currently, DynEx is built for individuals, but in the working world, designers and engineers usually work in teams. Traditionally, designers and engineers are siloed during software development - designers design and engineers implement, with product managers attempting to bridge the gap between the two. Context is bound to be lost when requirements are handed off across teams. Our user study with individuals demonstrated the value in contextualizing implementations in design, and these insights can be scaled to a team environment. The Design Matrix can bridge the gap between teams by providing a continuous feedback loop between designers and engineers. Design and engineering teams can contribute their unique perspectives and expertise in the matrix: designers can populate the Person through user interviews, backend engineers can define the Approach, and front-end engineers can describe the Interaction. This shared structure ensures that all perspectives are synthesized and valued, reducing inefficiencies from conflicting opinions and redundant meetings that often lack clear, actionable outcomes. It also offers a clear reference point for organizations to understand the relationship between design and code. Ultimately, the Design Matrix offers a more collaborative development process, bridging silos and providing context across teams.
@@ 8. Limitations and Future Work
Our user study was limited to ten participants, all of whom had computer science backgrounds and many years of technical experience. This may not be representative of the broader population that could benefit from an exploratory programming system — namely, anyone with an idea to prototype, regardless of their past technical exposure. Future studies should expand the demographics of participants to best inform the development of a system for users of varying technical backgrounds. Our study was also limited in the amount of time participants had to prototype. They had 30 minutes with each system — with additional time, participants may have created a variety of prototypes and compared them with one another to emulate a richer prototyping experience. Future studies should examine how users interact with the system when given more time.
Our user study also only used Claude Artifact as a baseline. Although we conducted a brief exploration which suggested Claude Artifact as being the most suitable as a baseline compared to other existing LLM-based tools, future work could involve comparisons between more systems to identify specific components or design choices that are effective in supporting exploratory programming. Furthermore, while users were given an introduction to Claude Artifact, they were still quite inexperienced; experienced users might use Claude Artifact as a tool for prototyping differently than novices.
@@ 9. Conclusion
The advancement of LLMs poses a unique opportunity for exploratory programming. LLMs can generate code rapidly and also help users enhance and refine their problem space. In this paper, we presented DynEx, a system to accelerate exploratory programming through a structured, LLM-guided Design Matrix and code-generation through Modular Stepwise Implementation. Through our user evaluation with 10 technical experts, we found that DynEx’s Design Matrix allowed to users to explore, refine, and realize feature-rich applications suitable for prototyping. We conclude by discussing how systems like DynEx can help bridge the gap between design and implementation, empowering a broad range of individuals to bring their ideas to life.
@@ Appendix A Few-shot Examples for the Design Matrix
OKCupid
PersonXIdea
Single person
PersonXGrounding
- Difficulty in finding potential partners who meet specific criteria like race, religion, age, or occupation.
- Challenges in efficiently filtering through dating apps to locate compatible matches.
- Need for an application that streamlines the process by allowing users to set precise criteria and receive curated suggestions.
ApproachXIdea
Searchable Database to allow people to search for people based on specified criteria
ApproachXGrounding
- Ensure the database includes comprehensive filters such as age, gender, sex, religion, and occupation to meet users’ specific search criteria.
- Develop a robust and scalable search algorithm that efficiently handles large datasets and returns accurate results based on the selected filters.
- Implement user-friendly search and filtering interfaces that make it easy for users to apply multiple criteria and refine their search results.
InteractionXIdea
Faceted Browsing
InteractionXGrounding
- Provide users with multiple facet filters, including age, gender, religion, occupation, and location, to refine their search effectively.
- Ensure the UI dynamically updates search results in real-time as users adjust their facet filters, offering immediate feedback.
- Design intuitive navigation with clear options to reset filters, switch criteria, and save searches, using checkboxes, sliders, and dropdowns for ease of use.
Tinder
PersonXIdea
Single person PersonXGrounding
- Users often struggle to find matches that meet their physical preferences quickly and efficiently.
- The abundance of profiles makes it challenging to identify those that align with specific looks or appearances.
- A streamlined approach to swiping and filtering by appearance would help users connect with potential matches faster, focusing on visual attraction. ApproachXIdea
Lower the cognitive load by providing less information, making it easier to judge potential matches quickly ApproachXGrounding
- Limit the displayed information to essential details, such as a single profile photo and a brief tagline, to encourage snap judgments.
- Focus on visual appeal as the primary matching criterion, reducing the need for users to sift through extensive profiles.
- Use an algorithm to prioritize matches based on visual preferences and minimal data inputs, streamlining the matching process. InteractionXIdea
Card Swipe InteractionXGrounding
- Each card should prominently feature a large profile photo, as visual appeal is the primary factor in this interaction.
- Include minimal text, such as the person’s name, age, and a short tagline or fun fact, to provide just enough context without overwhelming the user.
- Add simple icons or buttons for actions like ”Like” or ”Pass,” ensuring that users can quickly swipe or tap to make their choice.
Coffee Meets Bagel
PersonXIdea
Single person
PersonXGrounding
- Users who are looking for serious relationships prefer fewer, high-quality matches over endless swiping.
- The overwhelming number of potential matches on other apps can make it difficult to focus on finding a meaningful connection.
- An app designed for serious dating should streamline the process by offering a curated selection of potential partners, reducing time spent on the app. ApproachXIdea: Lower the cognitive load by having less matches to make more intentional judgements
ApproachXGrounding
- Present a select number of potential matches each day to prevent decision fatigue and promote thoughtful consideration.
- Display key information like shared interests, compatibility indicators, and mutual friends to aid decision-making without overwhelming the user.
- Prioritize quality over quantity, ensuring that each match is relevant to the user’s preferences and relationship goals.
InteractionXIdea
Feed with 5 options to date InteractionXGrounding
- The daily message should include a concise profile summary for each of the five matches, highlighting essential details such as name, age, occupation, and a short personal note or shared interest.
- Include compatibility scores or commonalities (e.g., mutual friends, hobbies) to help users quickly assess each match’s potential.
- Provide clear action buttons within the message to either like, pass, or start a conversation, making it easy for users to engage with their daily options.
Appendix B Few-shot Examples for the Project Specification
Music Recommendation App
Application Layout:
- Create a clean, simple interface divided into two main sections: ”Discover” and ”Favorites.”
- The ”Discover” section would have a large area that changes dynamically to display song details with each swipe (song name, artist, album, genre, and description).
- It would also have ’like’, ’dislike’ and ’skip’ buttons, which will work with simple clicks.
- The ’Favorites’ section would be a list of all liked songs.
User Interactions:
- The user can click to swipe a song left (dislike), right (like), or down (skip).
- The user can click on a song to save to favorites once liked. - The user can navigate between ”Discover” and ”Favorites” sections via top navigation tabs.
Inputs and Logic:
- The app will use the user’s interactions (likes, dislikes, skips) as inputs to an ML model (implemented using GPT) to evolve its music recommendations in real-time.
- On initial use, ask the user for favorite genres, artists, or songs to kickstart the ML algorithm.
- The user’s interaction with each song (whether they like, dislike, or skip it) will further tailor the recommendations.
- The saved favorite songs will be stored
- There is no need to create placeholder data for music, as GPT will return the music recommendations.
Outfit Generator App
Application Layout:
- Create a clean, minimalist interface with a prominent central area for displaying outfit recommendations.
- Divide the interface into three main sections: ”Outfit Recommendations,” ”Wardrobe,” and ”Saved Outfits.”
- The ”Outfit Recommendations” section should display swipeable cards with visual representations of the recommended outfits, along with relevant tags (season, occasion, style).
- The ”Wardrobe” section should allow users to input their clothing items, categorized by type (tops, bottoms, dresses, etc.).
- The ”Saved Outfits” section should display a grid of liked outfits for future reference.
User Interactions:
- Users can swipe left by clicking no, or right by clicking yes on the outfit recommendation cards to dislike or like the outfit, respectively.
- Users can click on individual clothing items in the ”Wardrobe” section to add or remove them from their virtual wardrobe.
- Users can click on a liked outfit in the ”Saved Outfits” section to view its details or remove it from the saved list.
Inputs and Logic:
- The app will use the user’s initial wardrobe inputs and style preferences (gathered through a brief questionnaire) to kickstart the GPT-powered outfit recommendation algorithm.
- The algorithm will consider factors like season, occasion, and the user’s wardrobe items to generate outfit recommendations.
- The user’s interactions (likes, dislikes) with the recommended outfits will be used as feedback to refine and personalize the algorithm’s recommendations over time.
- The liked outfits will be saved in the ”Saved Outfits” section for future reference.
- Create placeholder data for the initial wardrobe.
Plant Watering Tracker
Application Layout:
- Create a clean and intuitive interface with a prominent section for the ”Watering Calendar.”
- Divide the interface into three main sections: ”Watering Calendar,” ”Plant List,” and ”Watering Reminders.”
- The ”Watering Calendar” section should display a monthly calendar view with visual indicators for scheduled watering days.
- The ”Plant List” section should allow users to add and manage their plant collection, including details like species, pot size, and watering requirements.
- The ”Watering Reminders” section should display upcoming watering tasks and allow users to set notification preferences.
User Interactions:
- Users can click on specific dates in the ”Watering Calendar” to schedule or modify watering tasks for individual plants or groups.
- Users can click on plants in the ”Plant List” to view or edit their details, including watering schedules.
- Users can set notification preferences (email, push notifications, etc.) for upcoming watering tasks in the ”Watering Reminders” section.
Inputs and Logic:
- The app will use the user’s initial plant inputs (species, pot size, etc.) to determine baseline watering requirements for each plant.
- An algorithm (implemented using GPT) will analyze factors like plant type, pot size, and environmental conditions (temperature, humidity, etc.) to generate adaptive watering schedules.
- The algorithm will learn from the user’s interactions (manually adjusting watering schedules, plant health feedback) to refine its recommendations over time.
- The app will send reminders based on the user’s scheduled watering tasks and notification preferences.
- Create placeholder data for the user’s current plants.
Appendix C Prompt and Few-shot Examples for Placeholder Data
You are generating fake JSON data for a UI that a user wants to create. The spec should give instructions as to what data needs to be generated.
For example, for this spec:
”Application Layout:
- Create a clean, minimalist interface with a prominent central area for displaying outfit recommendations.
- Divide the interface into three main sections: ”Outfit Recommendations,” ”Wardrobe,” and ”Saved Outfits.”
- The ”Outfit Recommendations” section should display swipeable cards with visual representations of the recommended outfits, along with relevant tags (season, occasion, style).
- The ”Wardrobe” section should allow users to input their clothing items, categorized by type (tops, bottoms, dresses, etc.).
- The ”Saved Outfits” section should display a grid of liked outfits for future reference.
User Interactions:
- Users can swipe left by clicking no, or right by clicking yes on the outfit recommendation cards to dislike or like the outfit, respectively.
- Users can click on individual clothing items in the ”Wardrobe” section to add or remove them from their virtual wardrobe.
- Users can click on a liked outfit in the ”Saved Outfits” section to view its details or remove it from the saved list.
Inputs and Logic:
- The app will use the user’s initial wardrobe inputs and style preferences (gathered through a brief questionnaire) to kickstart the GPT-powered outfit recommendation algorithm.
- The algorithm will consider factors like season, occasion, and the user’s wardrobe items to generate outfit recommendations.
- The user’s interactions (likes, dislikes) with the recommended outfits will be used as feedback to refine and personalize the algorithm’s recommendations over time.
- The liked outfits will be saved in the ”Saved Outfits” section for future reference.
- Create placeholder data for initial wardrobe.” we only want to generate fake data for the INITIAL wardrobe, not the outfit recommendations.
Also, consider the user in this situation, and generate data tailored to the user if necessary. The application is for person_idea, with these details: person_grounding
Please generate a JSON array of fake data with appropriate fields. Here is an example:
Input: I want to create a UI that visualizes a beauty store’s inventory… It should have the fields ‘title‘, ‘description‘, ‘price‘, ‘discountPercentage‘, ‘rating‘, ‘stock‘, ‘brand‘, ‘category‘
System result:
                "id": 11,
                "title": "perfume Oil",
                "description": "Mega Discount",
                "price": 13,
                "discountPercentage": 8.4,
                "rating": 4.26,
                "stock": 65,
                "brand": "Impression of Acqua Di Gio",
                "category": "fragrances",
                "id": 12,
                "title": "perfume Oil",
                "description": "Half Off",
                "price": 15,
                "discountPercentage": 12.3,
                "rating": 3.46,
                "stock": 2343,
                "brand": "Victoria Secret",
                "category": "fragrances",
Please follow these rules while creating the JSON array
1. Please only return the JSON array and nothing else.
2. Array length should be length 10-20.
3. Please ensure that the generated data makes sense.
Appendix D Few-shot Examples for a Stepwise Plan Given Spec
Spec
Application Layout
- Create a clean, minimalist interface with a prominent central area for displaying outfit recommendations.
- Divide the interface into three main sections: ”Outfit Recommendations,” ”Wardrobe,” and ”Saved Outfits.”
- The ”Outfit Recommendations” section should display swipeable cards with visual representations of the recommended outfits, along with relevant tags (season, occasion, style).
- The ”Wardrobe” section should allow users to input their clothing items, categorized by type (tops, bottoms, dresses, etc.).
- The ”Saved Outfits” section should display a grid of liked outfits for future reference.
User Interactions:
- Users can swipe left by clicking no, or right by clicking yes on the outfit recommendation cards to dislike or like the outfit, respectively.
- Users can click on individual clothing items in the ”Wardrobe” section to add or remove them from their virtual wardrobe.
- Users can click on a liked outfit in the ”Saved Outfits” section to view its details or remove it from the saved list.
Inputs and Logic:
- The app will use the user’s initial wardrobe inputs and style preferences (gathered through a brief questionnaire) to kickstart the GPT-powered outfit recommendation algorithm.
- The algorithm will consider factors like season, occasion, and the user’s wardrobe items to generate outfit recommendations.
- The user’s interactions (likes, dislikes) with the recommended outfits will be used as feedback to refine and personalize the algorithm’s recommendations over time.
- The liked outfits will be saved in the ”Saved Outfits” section for future reference.
- Create placeholder data for the initial wardrobe.
Plan
1. Set up the React application and create the main layout with the three sections: ’Outfit Recommendations’, ’Wardrobe’, and ’Saved Outfits’. Read in the placeholder data from the endpoint.
2. Implement the ’Outfit Recommendations’ section with swipeable cards using MUI components. Create placeholder data for initial outfit recommendations.
3. Implement the ’Wardrobe’ section with a list of clothing items categorized by type (tops, bottoms, dresses, etc.). Allow users to add or remove items from their virtual wardrobe.
4. Implement the ’Saved Outfits’ section with a grid layout to display liked outfits. Allow users to view outfit details or remove outfits from the saved list.
5. Integrate GPT to generate outfit recommendations based on the user’s wardrobe and style preferences. Implement the logic to handle user interactions (likes, dislikes) and refine the recommendations accordingly.
Appendix E Prompt for a Basic Code Generation
The entire app will be in one index.html file. It will be written entirely in HTML, Javascript, and CSS. The design should not incorporate routes. Everything should exist within one page. No need for design mockups, wireframes, or external dependencies.
The entire app will be written using React and MUI. Load MUI from the CDN. Here is an example:
    <!DOCTYPE html>
<html lang="en">
<head>
  <meta charset="UTF-8">
  <meta name="viewport" content="width=device-width,
  initial-scale=1.0">
  <title>React App with MUI and Hooks</title>
  <!-- Load React and ReactDOM from CDN -->
  <script src="https://unpkg.com/react@18/umd/
  react.development.js" crossorigin></script>
  <script src="https://unpkg.com/react-dom@18/
  umd/react-dom.development.js" crossorigin></script>
  <!-- Babel for JSX transformation -->
  <script src="https://unpkg.com/@babel/standalone/
  babel.min.js"></script>
  <!-- Load MUI from CDN -->
  <link rel="stylesheet" href="https://fonts.googleapis
  .com/css?family=Roboto:300,400,500,700&display=swap" />
  <script src="https://unpkg.com/@mui/material@
  5.0.0-rc.1/umd/material-ui.development.js" crossorigin>
  </script>
</head>
<body>
  <div id="root"></div>
  <script type="text/babel">
    const {
      Button,
      Container,
      Typography,
      TextField,
    } = MaterialUI;
    const { useState, useEffect } = React;
    function App() {
      const [count, setCount] = useState(0);
      const [name, setName] = useState(’’);
      useEffect(() => {
        document.title = \‘Count: \\${count}\‘;
      }, [count]);
      return (
        <Container>
          <Typography variant="h2" component="h1"
          gutterBottom>
            Hello, React with Material-UI and Hooks!
          </Typography>
          <Typography variant="h5">
            Count: {count}
          </Typography>
          <Button variant="contained" color="primary"
          onClick={() => setCount(count + 1)}>
            Increment
          </Button>
          <TextField
            label="Name"
            value={name}
            onChange={(e) => setName(e.target.value)}
            variant="outlined"
            margin="normal"
            fullWidth
          <Typography variant="h6">
            Name: {name}
          </Typography>
        </Container>
    const rootElement = document.getElementById(’root’);
    const root = ReactDOM.createRoot(rootElement);
    root.render(<App />);
  </script>
</body>
</html>
- DO NOT DELETE PREVIOUS CODE. DO NOT RETURN A CODE SNIPPET. RETURN THE ENTIRE CODE. Only ADD to existing code to implement the task properly. DO NOT COMMENT PARTS OF THE CODE OUT AND WRITE /*…rest of the code */ or something similar. DO NOT COMMENT ANY PARTS OF THE CODE OUT. DO NOT COMMENT ANY PARTS OF THE CODE OUT FROM PREVIOUS CODES.
- DO NOT COMMENT /* Other sections remain the same */. ALL THE CODE MUST EXIST. ALL YOU ARE DOING IS ADDING FUNCTIONALITY. ADDING FUNCTIONALITY - DO NOT REMOVE ANY PREVIOUS CODE.
- The entire file should be less than 420 lines of code. MUST BE LESS THAN 420 LINES OF CODE.
- DO NOT LOAD ANYTHING ELSE IN THE CDN. Specifically, DO NOT USE: MaterialUI Icon, Material UI Lab.
- Do not return separate code files. All the components should be in one code file and returned.
- Do not type import statements. Assume that MUI and react are already imported libraries, so to use the components simply do so like this: const {Button, Container, Typography, TextField } = MaterialUI; or const { useState, useEffect } = React;
- DO NOT USE THESE MUI COMPONENTS: Calendar, DatePicker, TimePicker, SwipeableViews, SwipeableViewsVirtualizer, Fade, MopbileStepper as they do not exist.
Appendix F Few-shot Example for Self-Invoking GPT
try {{
 const response = await fetch(’https://api.openai.com/v1
 /chat/completions’, {{
   method: ’POST’,
   headers: {{
     ’Content-Type’: ’application/json’,
     ’Authorization’: ‘Bearer {openai_api_key}‘
   body: JSON.stringify({{
     model: ’gpt-4’,
     messages: [
         role: ’system’,
         content: ’You are a helpful assistant providing
         clothing recommendations based on user preferences.’
         role: ’user’,
         content: ‘
         Based on the following preferences, provide a list
         of recommended clothing items in the following JSON
         format:
             "id": 6,
             "itemName": "Striped T-Shirt",
             "description": "A classic striped t-shirt made
             of 100% cotton.",
             "imageUrl": "https://example.com/striped-
             tshirt.jpg",
             "size": "M",
             "brand": "H&M",
             "price": 19.99
         Preferences: XYX
         Ensure the JSON is valid and adheres strictly to
         this format. Do not type any additional text,
         only provide the JSON.‘
 const result = await response.json();
 const parsedData = JSON.parse(result.choices[0].message.
 content);
 setRecommendations(parsedData);
}} catch (err) {{
 setError(err.message);
}} finally {{
 setLoading(false);
Appendix G Few-shot Example for Self-Invoking OpenAI for Images
    try:
   response = await fetch(’https://api.openai.com/v1/chat
   /completions’, {{
       ’method’: ’POST’,
       ’headers’: {{
           ’Content-Type’: ’application/json’,
           ’Authorization’: ’Bearer {openai_api_key}’
       ’body’: json.dumps({{
           ’model’: ’gpt-4’,
           ’messages’: [
                   ’role’: ’system’,
                   ’content’: ’You are a helpful assistant
                   providing grocery recommendations based on
                   dietary preferences and restrictions.’
                   ’role’: ’user’,
                   ’content’: \"\"\"
                   Based on the following dietary preferences
                   and restrictions, provide a list of
                   recommended grocery items in the following
                   JSON format:
                           "name": "Organic Quinoa",
                           "description": "A gluten-free,
                           high-protein grain.",
                           "category": "Grains",
                           "nutritionalInfo": {{
                               "calories": 120,
                               "fat": 2.1,
                               "protein": 4.4,
                               "carbs": 21.3
                           "compatibility": "Vegan, Gluten-Free,
                           Vegetarian"
                   Preferences and Restrictions: {{preferences}}
                   Ensure the JSON is valid and adheres strictly
                   to this format. Do not type any additional
                   text, only provide the JSON.
   result = await response.json()
   parsed_data = json.loads(result[’choices’][0][’message’]
   [’content’])
   # Fetch images for each recommendation
   for item in parsed_data:
       image_response = await fetch(’https://api.openai.com/
       v1/images/generations’, {{
           ’method’: ’POST’,
           ’headers’: {{
               ’Content-Type’: ’application/json’,
               ’Authorization’: ’Bearer {openai_api_key}’
           ’body’: json.dumps({{
               ’prompt’: f"An image of {{item[’name’]}},
               a {{item[’category’]}} item.",
               ’n’: 1,
               ’size’: "256x256"
       image_result = await image_response.json()
       item[’imageUrl’] = image_result[’data’][0][’url’]
   set_recommendations(parsed_data)
except Exception as err:
   print(err)
   set_error(str(err))
finally:
   set_loading(False)
Appendix H Prompt for Iteration Within Steps to Debug
User Message
Please fix the problem that the user describes: {problem}
System Message
A coding task has been implemented for a project we are working on. For context, this is the project description:
{spec}. The task was this: {task }. This is the faked_data: {faked_data}. However, the task was not implemented fully correctly. The user explains what is wrong in the problem {problem}. There is already existing code in the index.html file. Using the existing code {task_code}. Please fix the problem. PLEASE DO NOT DELETE EXISTING CODE. ONLY FIX THE BUG. Return the FULL CODE NEEDED TO HAVE THE APP WORK, INSIDE THE INDEX.HTML file.
Appendix I Design Matrix for Usage Scenario
Person:Idea
1. Non-native speakers interested in Chinese culture 2. Visual learners struggling with language memorization 3. Travel enthusiasts planning a trip to China 4. Retired adult wanting to expand linguistic skills 5. Busy university student wanting to study Chinese in his free time 6. Travel enthusiasts planning a trip to China
Person:Grounding
- Confusion arises due to unfamiliarity with Chinese characters and their complex structure, slowing the learning process. - Difficulty in linking characters to their corresponding meaning or pronunciation, hindering vocabulary acquisition. - Traditional memorization methods offer little aid to visual learners who could better recall information through imagery.
Approach:Idea (1)
1. Pictorial spaced repetition learning 2. Visual storytelling for language acquisition 3. Cognitive load theory for efficient memorization
Approach:Grounding (1)
- Implement the spaced repetition system (SRS) algorithm to schedule review times according to each user’s progress, allowing items to reappear before they’re likely to be forgotten. - Incorporate visual images representing Chinese characters or words, enhancing the memory association between visual cues and corresponding meanings. - Integrate regular reviews into the learning process, which are fine-tuned according to the user’s performance, to further solidify memory and retention. - Leverage cognitive science principles such as interleaving (mixing similar tasks) and retrieval practice (recalling an item from memory), to increase learning effectiveness.
Approach:Idea (2)
1. Pictorial spaced repetition learning 2. Visual storytelling for language acquisition 3. Cognitive load theory for efficient memorization
Approach:Grounding (2)
- Integrate visual aids, such as illustrations or animations, that correlate with each word’s meaning to enhance understanding and recall. - Develop an algorithm that links related words and images together in a meaningful story, promoting stronger memory associations. - Make use of GPT to generate context-rich sentences or mini-stories, helping to create a narrative around each word or character. - Ensure design of the learning material caters to visual learners with a focus on vibrant, engaging, and contextually relevant graphical representations.
Interaction:Idea
1. Simple guess-and-review quiz interface 2. Visual dictionary flashcard interface 3. Image-based language learning game interface
Interaction:Grounding
- The quiz interface should present the Chinese character or word along with its corresponding image. - The user then attempts to guess its meaning. If they respond accurately, the item is pushed back into the review cycle based on the SRS algorithm. - If the guess is incorrect, the correct meaning is displayed, and the item is scheduled for another review sooner. - Users should have a clear view of their progress and a way to navigate to previously learned words for self-study.
Appendix J Spec for Usage Scenario
Application Layout:
- Divide the interface into three main sections: ”Flashcards,” ”Quiz,” and ”Progress.”
- The ”Flashcards” section displays the Chinese character/word along with its corresponding image for visual association.
- The ”Quiz” section presents the character/word and image, prompting the user to input the meaning.
- The ”Progress” section shows the user’s performance metrics, such as accuracy rate and words mastered.
User Interactions:
- In the ”Flashcards” section, users can click through the flashcards to study the character/word and its associated image.
- In the ”Quiz” section, users can type in their answer for the displayed character/word and image.
- Users can click a ”Show Answer” button to reveal the correct meaning if they are unsure.
- Users can navigate between the ”Flashcards,” ”Quiz,” and ”Progress” sections using tabs or buttons.
Inputs and Logic:
- The app will use a pre-generated dataset of Chinese characters/words, their meanings, and associated images.
- The spaced repetition system (SRS) algorithm will determine the optimal interval for reviewing each character/word based on the user’s performance.
- Correctly answered items in the ”Quiz” section will have their review interval increased (pushed further back in the review cycle).
- Incorrectly answered items will have their review interval decreased (scheduled for more frequent review).
- The app will track the user’s progress, such as the number of words mastered and accuracy rates, and display this information in the ”Progress” section.
- Incorporate cognitive science principles like interleaving (mixing different types of content) and retrieval practice (actively recalling information) into the ”Quiz” section to enhance learning effectiveness.
By combining visual associations with images, spaced repetition, and interactive quizzes, this UI aims to provide an engaging and effective way for non-Chinese speaking students to learn and retain Chinese characters and vocabulary.
Appendix K Placeholder Data for Usage Scenario
{CJK}
UTF8gbsn
    "id": 1,
    "chinese": "",
    "meaning": "I study Chinese",
    "imagePath": "images/studying.jpg"
    "id": 2,
    "chinese": "?",
    "meaning": "How are you?",
    "imagePath": "images/greeting.jpg"
    "id": 3,
    "chinese": "",
    "meaning": "This is a book",
    "imagePath": "images/book.jpg"
    "id": 4,
    "chinese": "",
    "meaning": "I like to eat apples",
    "imagePath": "images/apple.jpg"
    "id": 5,
    "chinese": "",
    "meaning": "The weather is nice today",
    "imagePath": "images/sunny.jpg"
    "id": 6,
    "chinese": "",
    "meaning": "Let’s go to the park",
    "imagePath": "images/park.jpg"
    "id": 7,
    "chinese": "",
    "meaning": "This room is big",
    "imagePath": "images/room.jpg"
    "id": 8,
    "chinese": "",
    "meaning": "I like to listen to music",
    "imagePath": "images/music.jpg"
    "id": 9,
    "chinese": "",
    "meaning": "Let’s go watch a movie",
    "imagePath": "images/movie.jpg"
    "id": 10,
    "chinese": "",
    "meaning": "I want to buy a new piece of clothing",
    "imagePath": "images/shopping.jpg"
    "id": 11,
    "chinese": "",
    "meaning": "Let’s go eat dinner",
    "imagePath": "images/dinner.jpg"
    "id": 12,
    "chinese": "",
    "meaning": "This city is bustling",
    "imagePath": "images/city.jpg"
    "id": 13,
    "chinese": "",
    "meaning": "I like to travel",
    "imagePath": "images/travel.jpg"
    "id": 14,
    "chinese": "",
    "meaning": "Let’s go exercise",
    "imagePath": "images/exercise.jpg"
    "id": 15,
    "chinese": "",
    "meaning": "This project is interesting",
    "imagePath": "images/project.jpg"
Appendix L Plan for Usage Scenario
  1. (1)
    Set up the React application and create the main layout with the three sections: ’Flashcards’, ’Quiz’, and ’Progress’. Read in the pre-generated dataset of Chinese characters/words, their meanings, and associated images from the data endpoint.
  2. (2)
    Implement the ’Flashcards’ section with a display for the Chinese character/word and its corresponding image. Allow users to click through the flashcards.
  3. (3)
    Implement the ’Quiz’ section with a display for the Chinese character/word and its associated image. Allow users to input their answer and reveal the correct meaning. Integrate the spaced repetition system (SRS) algorithm to determine the optimal review interval for each character/word based on the user’s performance.
  4. (4)
    Implement the ’Progress’ section to display the user’s performance metrics, such as accuracy rate and words mastered. Incorporate cognitive science principles like interleaving and retrieval practice into the ’Quiz’ section.
  5. (5)
    Call GPT to generate images for the Chinese characters/words and associate them with the corresponding entries in the dataset. Display these images in the ’Flashcards’ and ’Quiz’ sections.